OpenStudy (anonymous):

7 – 4x = 31 A. x = 6 B. x = 9.5 C. x = –6 D. x = –9.5

OpenStudy (anonymous):

Subtract 7 from 31. then divide by 4

OpenStudy (anonymous):

7-4x=31 -4x=24 x=-6

OpenStudy (anonymous):

@study100 dont just give the answer please we are not here to do their homework for them

OpenStudy (anonymous):

this question is hard

OpenStudy (anonymous):

 Basically you are trying to get x on one side and the constants on the other then you must have x without a coefficient and when you do that. Wallah an answer :D